Please help me with this : f(x)=8x^2+6 and g(x)=6x+a, find a so that the graph of f of g crosses the y-axis at 78
Answer by stanbon(75887)

Please help me with this : f(x)=8x^2+6 and g(x)=6x+a, find a so that the graph of f of g crosses the y-axis at 78
----------
fog(x) = f[g(x)] = f[6x+a] = 8(6x+a)^2+6
= 8(36x^2+12ax+a^2)+6
= 288x^2 + 96ax + 8a^2+6
----
Solve for "a" based using (0,78)
----------
78 = 0 + 0 + 8a^2+6
8a^2 = 72
a^2 = 9
a = 3 or a = -3
====================
Cheers,
Stan H.
====================